# Zero-sum partitions of Abelian groups and their applications to magic- and antimagic-type labelings

### v4

@@ Abstract.
2010 Mathematics Subject Classification:
05E16, 20K01, 05C25, 05C78
This work was supported by program “Excellence initiative – research university” for the AGH University.
1. Introduction
1.1. Preliminaries
For standard terms and notation in graph theory, the reader is referred to the textbook by Diestel [10] and the monograph by Brandstädt et al. [3], an introduction to magic- and antimagic-type labelings can be found in the monograph by Bača et al. [2], and the fundamental results in abstract algebra that we use can be found in the textbook by Gallian [13].
@@ 1.1. Preliminaries
The fundamental theorem of finite Abelian groups states that every finite Abelian group $\Gamma$ is isomorphic to the direct product of some cyclic subgroups of prime-power orders [13]. In other words, there exists a positive integer $k$, (not necessarily distinct) prime numbers $\{p_{i}\}_{i=1}^{k}$, and positive integers $\{\alpha_{i}\}_{i=1}^{k}$, such that
@@
Recall that the sum of all elements of a group $\Gamma$ is equal to the sum of its involutions and the identity element. The following lemma is well known. The readers can consult [9] for a proof.
Lemma 1.1 ([9]).
@@ 1.2. Main Problem
Note that, by Lemma 1.1, an Abelian group $\Gamma$ with $|I(\Gamma)|=1$ does not admit zero-sum partitions (for every $t$ and $\{m_{i}\}_{i=1}^{t}$). Moreover, if $\{m_{i}\}_{i=1}^{t}$ is realizable in $\Gamma^{*}$, then necessarily $m_{i}\geq 2$ for every $1\leq i\leq t$. It was proved that this condition is sufficient if and only if $|I(\Gamma)|=0$ [22, 24] or $|I(\Gamma)|=3$ [24]. I.e., in these cases, for every positive integer $t$ and integer partition $\{m_{i}\}_{i=1}^{t}$ of $m-1$ with $m_{i}\geq 2$ for every $1\leq i\leq t$, $\{m_{i}\}_{i=1}^{t}$ is realizable in $\Gamma^{*}$ (with $\{w_{i}\}_{i=1}^{t}$ where $w_{i}=0$ for every $1\leq i\leq t$). In [7], we generalized this condition with the following definition.
@@
In 1957, inspired by Steiner triples research, Skolem posed the following question [20]: For $n\equiv 1\pmod{6}$, does there exist a partition of the set of non-zero elements of the cyclic group $\mathbb{Z}_{n}$ into triples such that the sum of elements in each subset is congruent to $0\pmod{n}$?
This question received an affirmative answer [15, 20, 21] and served as a starting point for a more general problem posed by Tannenbaum. Namely, we call a $6$-subset $C$ of an Abelian group $\Gamma$ good if $C=\{c,d,-c-d,-c,-d,c+d\}$ for some $c$ and $d$ in $\Gamma$. Notice that the sum of elements of a good $6$-subset is 00. Moreover, it can be partitioned into three zero-sum $2$-subsets or two zero-sum $3$-subsets.
The following definition was given by Tannenbaum [22].
@@
The following theorem was first conjectured by Kaplan et al. [16] (they also showed the necessity), and later proved by Zeng [24].
Theorem 1.6 ([24]).
@@
Note that, in general, even the weaker version of Conjecture 1.7 posed in [6] is still open.
Conjecture 1.9 ([6]).
@@
For every Abelian group $\Gamma$ with more than one involution and large enough order, first Cichacz and Tuza [8] showed that it has $4$-ZSPP, and later Müyesser and Pokrovskiy showed that it indeed has $3$-ZSPP [18]. This result was improved by Cichacz and Suchan [7], for every $\Gamma$ of order $2^{n}$ for some integer $n>1$ such that $I(\Gamma)\neq 1$, with the following theorem.
Theorem 1.10 ([7]).
@@
Let $\Gamma$ be an Abelian group such that $\Gamma\cong L\times H$, where $|L|=2^{\eta}$ and $|H|=\rho$ for a positive integer $\eta$ and an odd positive integer $\rho$, with $\rho>1$. The main contribution of this paper is to show that $\Gamma$ has $3$-ZSPP if $(\rho\bmod{6})\in\{1,3\}$. Moreover, for $(\rho\bmod{6})=5$, we show that $\Gamma$ has $4$-ZSPP. Therefore we improve the result from [8] and, towards proving Conjecture 1.9, we leave open only the question if, for $(\rho\bmod{6})=5$, $\Gamma$ has not only $4$-ZSPP, but also $3$-ZSPP. We also show that Conjecture 1.7 holds only for a group $\Gamma$ with $|I(\Gamma)|=3$ or $\Gamma\cong(\mathbb{Z}_{2})^{n}$ with any $n>1$. As a complement, we present some applications in irregular, magic-type, and antimagic-type graph labeling to illustrate the relations of zero-sum group partitioning to graph theory.
Our proofs use as base cases zero-sum partitions of $\left((\mathbb{Z}_{2})^{3}\times\mathbb{Z}_{3}\right)^{*}$, $\left((\mathbb{Z}_{2})^{3}\times\mathbb{Z}_{5}\right)^{*}$, and $\left((\mathbb{Z}_{2})^{3}\times\mathbb{Z}_{7}\right)^{*}$. They were analyzed by a computer program that we created, and sample zero-sum partitions that certify that the corresponding sequences are realizable are given in the annexes.
2. Main Result
Let Bij$(\Gamma)$ denote the set of all bijections from $\Gamma$ to itself. A complete mapping of a group $A$ is defined as $\varphi\in$Bij$(A)$ such that the mapping $\theta\colon g\mapsto g^{-1}\varphi(g)$ is also bijective (some authors refer to $\theta$, rather than $\varphi$, as the complete mapping) [17]. In the proof of our main result we will use the following property of complete mappings:
Lemma 2.1 ([6, 24]).
@@ 2. Main Result
$H=\{0\}\cup\left(\bigcup\limits_{i=1}^{|H^{*}|/6}\{b_{i},c_{i},-b_{i}-c_{i},-b_{i},-c_{i},b_{i}+c_{i}\}\right).$
@@
$\Gamma^{*}=\bigcup\limits_{a\in L^{*}}\{(a,0)\}\cup\left(\bigcup\limits_{a\in L}\bigcup\limits_{i=1}^{|H^{*}|/6}\{(a,b_{i}),(\phi(a),c_{i}),(\varphi(a),-b_{i}-c_{i}),(-a,-b_{i}),(-\phi(a),-c_{i}),(-\varphi(a),b_{i}+c_{i})\}\right),$
@@
$\cup\left(\bigcup\limits_{a\in(\mathbb{Z}_{2})^{3}\setminus\{i_{0},i_{1},i_{2},i_{3}\}}\{(a,b_{1}),(\phi(a),c_{1}),(\varphi(a),-b_{1}-c_{1}),(-a,-b_{1}),(-\phi(a),-c_{1}),(-\varphi(a),b_{1}+c_{1})\}\right)$
$\cup\left(\bigcup\limits_{a\in(\mathbb{Z}_{2})^{3}}\bigcup\limits_{i=2}^{|H^{*}|/6}\{(a,b_{i}),(\phi(a),c_{i}),(\varphi(a),-b_{i}-c_{i}),(-a,-b_{i}),(-\phi(a),-c_{i}),(-\varphi(a),b_{i}+c_{i})\}\right).$
@@
$\bigcup\limits_{a\in((\mathbb{Z}_{2})^{3})^{*}}\{(a,0)\}\cup\left(\bigcup\limits_{a\in(\mathbb{Z}_{2})^{3}}\{(a,b_{1}),(\phi(a),c_{1}),(\varphi(a),-b_{1}-c_{1}),(-a,-b_{1}),(-\phi(a),-c_{1}),(-\varphi(a),b_{1}+c_{1})\}\right)$
@@
We will need now the following lemma (see Lemma 2.4 in [7]):
Lemma 2.3 ([7]).
@@
$H=\{0,b,-b\}\cup\left(\bigcup\limits_{i=1}^{(|H^{*}|-2)/6}\{b_{i},c_{i},-b_{i}-c_{i},-b_{i},-c_{i},b_{i}+c_{i}\}\right).$
@@
$\Gamma^{*}=\{(0,b),(0,-b)\}\cup\left(\bigcup\limits_{a\in L^{*}}\{(a,0),(\phi(a),b),(\varphi(a),-b)\}\right)\cup W.$
@@
$\Gamma^{*}=M^{*}\cup\left(\bigcup\limits_{a\in L}\{(a,b),(-a,-b)\}\right)\cup W.$
@@
$\Gamma^{*}=M^{*}\cup\left(\bigcup\limits_{a\in L}\{(a,1),(-a,2)\}\right).$
@@
$\begin{array}[]{ll}\Gamma^{*}=&B^{*}\cup\left(\bigcup\limits_{\begin{subarray}{c}g\in B\end{subarray}}\{a+g,-a-g)\}\cup\bigcup\limits_{\begin{subarray}{c}g\in B\end{subarray}}\{e_{1}+g,e_{2}+\phi(g),-e_{1}-e_{2}+\varphi(g)\}\right)\\ &\cup\left(\bigcup\limits_{g\in B}\left\{c+g,d+\phi(g),-c-d+\varphi(g),-c-g,-d-\phi(g),c+d-\varphi(g)\right\}\right),\end{array}$
@@
$\left(\bigcup\limits_{g\in B}\left\{c+g,d+\phi(g),-c-d+\varphi(g),-c-g,-d-\phi(g),c+d-\varphi(g)\right\}\right).$
@@
$\begin{array}[]{ll}\Gamma^{*}=&\{a,-a\}\cup\left(\bigcup\limits_{\begin{subarray}{c}g\in B^{*}\end{subarray}}\{a+g,\phi(g),-a+\varphi(g)\}\cup\bigcup\limits_{\begin{subarray}{c}g\in B\end{subarray}}(\{e_{1}+g,e_{2}+\phi(g),-e_{1}-e_{2}+\varphi(g)\}\right)\\ &\cup\left(\bigcup\limits_{g\in B}\left\{c+g,d+\phi(g),-c-d+\varphi(g),-c-g,-d-\phi(g),c+d-\varphi(g)\right\}\right).\end{array}$
@@
$H=\{0,b,-b,c,-c\}\cup\left(\bigcup\limits_{i=1}^{(|H^{*}|-4)/6}\{b_{i},c_{i},-b_{i}-c_{i},-b_{i},-c_{i},b_{i}+c_{i}\}\right).$
@@
$\Gamma^{*}=\{(0,b),(0,-b)\}\cup\left(\bigcup\limits_{a\in L^{*}}\{(a,0),(\phi(a),b),(\varphi(a),-b)\}\right)\cup\left(\bigcup\limits_{a\in L}\{(a,c),(-a,-c)\}\right)\cup W.$
@@
$\Gamma^{*}=M^{*}\cup\left(\bigcup\limits_{a\in L}\{(a,b),(-a,-b)\}\right)\cup\left(\bigcup\limits_{a\in L}\{(a,c),(-a,-c)\}\right)\cup W.$
@@
$\Gamma^{*}=M^{*}\cup\left(\bigcup\limits_{a\in L}\{(a,b),(-a,-b)\}\right)\cup\left(\bigcup\limits_{a\in L}\{(a,c),(-a,-c)\}\right).$
@@
$\begin{array}[]{ll}\Gamma^{*}=&B^{*}\cup\left(\bigcup\limits_{\begin{subarray}{c}g\in B\end{subarray}}\{a+g,-a-g\}\right)\cup\left(\bigcup\limits_{\begin{subarray}{c}g\in B\end{subarray}}\{b+g,-b-g\}\right)\\ &\cup\left(\bigcup\limits_{\begin{subarray}{c}g\in B\end{subarray}}\{e_{1}+g,e_{2}+\phi(g),-e_{1}-e_{2}+\varphi(g)\}\right)\\ &\cup\left(\bigcup\limits_{\begin{subarray}{c}g\in B\\ i=1,2\end{subarray}}\left\{c_{i}+g,d_{i}+\phi(g),-c_{i}-d_{i}+\varphi(g),-c_{i}-g,-d_{i}-\phi(g),c_{i}+d_{i}-\varphi(g)\right\}\right),\end{array}$
@@
$\left(\bigcup\limits_{g\in B}\left\{c_{1}+g,d_{1}+\phi(g),-c_{1}-d_{1}+\varphi(g),-c_{1}-g,-d_{1}-\phi(g),c_{1}+d_{1}-\varphi(g)\right\}\right).$
@@
$\begin{array}[]{ll}\Gamma^{*}=&\{a,-a\}\cup\left(\bigcup\limits_{\begin{subarray}{c}g\in B^{*}\end{subarray}}\{a+g,\phi(g),-a+\varphi(g)\}\right)\cup\left(\bigcup\limits_{\begin{subarray}{c}g\in B\end{subarray}}\{b+g,-b-g\}\right)\\ &\cup\left(\bigcup\limits_{\begin{subarray}{c}g\in B\end{subarray}}\{e_{1}+g,e_{2}+\phi(g),-e_{1}-e_{2}+\varphi(g)\}\right)\\ &\cup\left(\bigcup\limits_{\begin{subarray}{c}g\in B\\ i=1,2\end{subarray}}\left\{c_{i}+g,d_{i}+\phi(g),-c_{i}-d_{i}+\varphi(g),-c_{i}-g,-d_{i}-\phi(g),c_{i}+d_{i}-\varphi(g)\right\}\right),\end{array}$
@@
Since $|B|=|\Gamma|/20=|L|/4$, the sequence $(|L|/2-1,0,0)$ is realizable in $\left(\cup_{g\in B^{*}}\{a+g,\phi(g),-a+\varphi(g)\}\right)$ $\cup\left(\cup_{g\in B}\{e_{1}+g,e_{2}+\phi(g),-e_{1}-e_{2}+\varphi(g)\}\right)$. Since $s<|L|$ and $10\leq s-|L|/2+$ 1, we can find the remaining $s-|L|/2+1\leq|L|/2=2|B|$ zero-sum $3$-subsets in
@@
Assume first that $e$ is even. Let $m_{1},\dots,m_{|R|/2-1}=2$, $m_{|R|/2},\ldots,m_{|R|/2+|I(\Gamma)|/3-3}=3$, $m_{|R|/2+|I(\Gamma)|/3-2},$ $m_{|R|/2+|I(\Gamma)|/3-1}=4$. Note that $\sum_{i=1}^{|R|/2+|I(\Gamma)|/3-1}m_{i}=|\Gamma^{*}|$. Suppose that there is a subset partition $\{S_{i}\}_{i=1}^{t}$ of $\Gamma^{*}$ such that $|S_{i}|=m_{i}$ and $\sum_{s\in S_{i}}s=0$ for all $i$. Then $R\setminus\left(\bigcup_{i=1}^{|R|/2-1}S_{i}\right)=\{a,-a\}$ for some $a\in R$. Thus $a\in S_{j}$ for some $j\in\{|R|/2,\ldots,|R|/2+|I(\Gamma)|/3-1\}$ but then $\sum_{s\in S_{j}}s\neq 0$ since $I(\Gamma)\cup\{0\}$ is a subgroup of $\Gamma$. Thus for any $\iota_{1},\iota_{2},\iota_{3}\in I(\Gamma)$ there is $\iota_{1}+\iota_{2}\in I(\Gamma)$, $\iota_{1}+\iota_{2}+\iota_{3}\in I(\Gamma)\cup\{0\}$ but $a\notin I(\Gamma)\cup\{0\}$, a contradiction.

### v5

@@ Abstract
1 Introduction
1.1 Preliminaries
For standard terms and notation in graph theory, the reader is referred to the textbook by Diestel (2017) and the monograph by Brandstädt et al. (1999), an introduction to magic- and antimagic-type labelings can be found in the monograph by Bača et al. (2019), and the fundamental results in abstract algebra that we use can be found in the textbook by Gallian (2016).
@@ 1.1 Preliminaries
The fundamental theorem of finite Abelian groups states that every finite Abelian group $\Gamma$ is isomorphic to the direct product of some cyclic subgroups of prime-power orders (Gallian, 2016). In other words, there exists a positive integer $k$, (not necessarily distinct) prime numbers $\{p_{i}\}_{i=1}^{k}$, and positive integers $\{\alpha_{i}\}_{i=1}^{k}$, such that
@@
Recall that the sum of all elements of a group $\Gamma$ is equal to the sum of its involutions and the identity element. The following lemma is well known. The readers can consult Combe et al. (2004) for a proof.
Lemma 1.1 ((Combe et al., 2004)).
@@ 1.2 Main Problem
Note that, by Lemma 1.1, an Abelian group $\Gamma$ with $|I(\Gamma)|=1$ does not admit zero-sum partitions (for every $t$ and $\{m_{i}\}_{i=1}^{t}$). Moreover, if $\{m_{i}\}_{i=1}^{t}$ is realizable in $\Gamma^{*}$, then necessarily $m_{i}\geq 2$ for every $1\leq i\leq t$. It was proved that this condition is sufficient if and only if $|I(\Gamma)|=0$ (Tannenbaum, 1981; Zeng, 2015) or $|I(\Gamma)|=3$ (Zeng, 2015). I.e., in these cases, for every positive integer $t$ and integer partition $\{m_{i}\}_{i=1}^{t}$ of $m-1$ with $m_{i}\geq 2$ for every $1\leq i\leq t$, $\{m_{i}\}_{i=1}^{t}$ is realizable in $\Gamma^{*}$ (with $\{w_{i}\}_{i=1}^{t}$ where $w_{i}=0$ for every $1\leq i\leq t$). In Cichacz and Suchan (2023), we generalized this condition with the following definition.
@@
In 1957, inspired by Steiner triples research, Skolem posed the following question (Skolem, 1957): For $n\equiv 1\pmod{6}$, does there exist a partition of the set of non-zero elements of the cyclic group $\mathbb{Z}_{n}$ into triples such that the sum of elements in each subset is congruent to $0\pmod{n}$?
This question received an affirmative answer (Hanani, 1960; Skolem, 1957, 1958) and served as a starting point for a more general problem posed by Tannenbaum. Namely, we call a $6$-subset $C$ of an Abelian group $\Gamma$ good if $C=\{c,d,-c-d,-c,-d,c+d\}$ for some $c$ and $d$ in $\Gamma$. Notice that the sum of elements of a good $6$-subset is 00. Moreover, it can be partitioned into three zero-sum $2$-subsets or two zero-sum $3$-subsets.
The following definition was given by Tannenbaum (1981).
@@
The following theorem was first conjectured by Kaplan et al. (2009) (they also showed the necessity), and later proved by Zeng (2015).
Theorem 1.6 ((Zeng, 2015)).
@@
Note that, in general, even the weaker version of Conjecture 1.7 posed by Cichacz (2018) is still open.
Conjecture 1.9 ((Cichacz, 2018)).
@@
For every Abelian group $\Gamma$ with more than one involution and large enough order, first Cichacz and Tuza (2022) showed that it has $4$-ZSPP, and later Müyesser and Pokrovskiy (2022) showed that it indeed has $3$-ZSPP. This result was improved by Cichacz and Suchan (2023), for every $\Gamma$ of order $2^{n}$ for some integer $n>1$ such that $I(\Gamma)\neq 1$, with the following theorem.
Theorem 1.10 ((Cichacz and Suchan, 2023)).
@@
Let $\Gamma$ be an Abelian group such that $\Gamma\cong L\times H$, where $|L|=2^{\eta}$ and $|H|=\rho$ for a positive integer $\eta$ and an odd positive integer $\rho$, with $\rho>1$. The main contribution of this paper is to show that $\Gamma$ has $3$-ZSPP if $(\rho\bmod{6})\in\{1,3\}$. Moreover, for $(\rho\bmod{6})=5$, we show that $\Gamma$ has $4$-ZSPP. Therefore we improve the result by Cichacz and Tuza (2022) and, towards proving Conjecture 1.9, we leave open only the question if, for $(\rho\bmod{6})=5$, $\Gamma$ has not only $4$-ZSPP, but also $3$-ZSPP. We also show that Conjecture 1.7 holds only for a group $\Gamma$ with $|I(\Gamma)|=3$ or $\Gamma\cong(\mathbb{Z}_{2})^{n}$ with any $n>1$. As a complement, we present some applications in irregular, magic-type, and antimagic-type graph labeling to illustrate the relations of zero-sum group partitioning to graph theory.
In proofs, we use zero-sum partitions of $\left((\mathbb{Z}_{2})^{3}\times\mathbb{Z}_{3}\right)^{*}$, $\left((\mathbb{Z}_{2})^{3}\times\mathbb{Z}_{5}\right)^{*}$, and $\left((\mathbb{Z}_{2})^{3}\times\mathbb{Z}_{7}\right)^{*}$ as base cases. They were analyzed by a computer program that we created, and sample zero-sum partitions that certify that the corresponding sequences are realizable are given in the annexes.
2 Main Result
Let Bij$(\Gamma)$ denote the set of all bijections from $\Gamma$ to itself. A complete mapping of a group $A$ is defined as $\varphi\in$Bij$(A)$ such that the mapping $\theta\colon g\mapsto g^{-1}\varphi(g)$ is also bijective (some authors refer to $\theta$, rather than $\varphi$, as the complete mapping) (Mann, 1944). In the proof of our main result we will use the following property of complete mappings:
Lemma 2.1 ((Cichacz, 2018; Zeng, 2015)).
@@ 2 Main Result
$H=\{0\}\cup\bigcup\limits_{i=1}^{|H^{*}|/6}\{b_{i},c_{i},-b_{i}-c_{i},-b_{i},-c_{i},b_{i}+c_{i}\}.$
@@
$\Gamma^{*}=\bigcup\limits_{a\in L^{*}}\{(a,0)\}$
$\cup\bigcup\limits_{a\in L}\bigcup\limits_{i=1}^{|H^{*}|/6}\{(a,b_{i}),(\phi(a),c_{i}),(\varphi(a),-b_{i}-c_{i}),(-a,-b_{i}),(-\phi(a),-c_{i}),(-\varphi(a),b_{i}+c_{i})\},$
@@
$\cup\bigcup\limits_{a\in(\mathbb{Z}_{2})^{3}\setminus\{i_{0},i_{1},i_{2},i_{3}\}}\{(a,b_{1}),(\phi(a),c_{1}),(\varphi(a),-b_{1}-c_{1}),(-a,-b_{1}),(-\phi(a),-c_{1}),(-\varphi(a),b_{1}+c_{1})\}$
$\cup\bigcup\limits_{a\in(\mathbb{Z}_{2})^{3}}\bigcup\limits_{i=2}^{|H^{*}|/6}\{(a,b_{i}),(\phi(a),c_{i}),(\varphi(a),-b_{i}-c_{i}),(-a,-b_{i}),(-\phi(a),-c_{i}),(-\varphi(a),b_{i}+c_{i})\}.$
@@
${\bigcup_{a\in((\mathbb{Z}_{2})^{3})^{*}}^{}}\{(a,0)\}\cup\smashoperator[]{\bigcup_{a\in(\mathbb{Z}_{2})^{3}}^{}}\{(a,b_{1}),(\phi(a),c_{1}),(\varphi(a),-b_{1}-c_{1}),(-a,-b_{1}),(-\phi(a),-c_{1}),(-\varphi(a),b_{1}+c_{1})\}$
@@
We will need now the following lemma (Cichacz and Suchan, 2023, Lemma 2.4):
Lemma 2.3 ((Cichacz and Suchan, 2023)).
@@
$H=\{0,b,-b\}\cup\bigcup\limits_{i=1}^{(|H^{*}|-2)/6}\{b_{i},c_{i},-b_{i}-c_{i},-b_{i},-c_{i},b_{i}+c_{i}\}.$
@@
$\Gamma^{*}=\{(0,b),(0,-b)\}\cup\bigcup\limits_{a\in L^{*}}\{(a,0),(\phi(a),b),(\varphi(a),-b)\}\cup W.$
@@
$\Gamma^{*}=M^{*}\cup\bigcup\limits_{a\in L}\{(a,b),(-a,-b)\}\cup W.$
@@
$\Gamma^{*}=M^{*}\cup\bigcup\limits_{a\in L}\{(a,1),(-a,2)\}.$
@@
$\begin{array}[]{ll}\Gamma^{*}=&B^{*}\cup\bigcup\limits_{\begin{subarray}{c}g\in B\end{subarray}}\{a+g,-a-g)\}\cup\bigcup\limits_{\begin{subarray}{c}g\in B\end{subarray}}\{e_{1}+g,e_{2}+\phi(g),-e_{1}-e_{2}+\varphi(g)\}\\ &\cup\bigcup\limits_{g\in B}\left\{c+g,d+\phi(g),-c-d+\varphi(g),-c-g,-d-\phi(g),c+d-\varphi(g)\right\},\end{array}$
@@
$\begin{array}[]{ll}\Gamma^{*}=&\{a,-a\}\cup\bigcup\limits_{\begin{subarray}{c}g\in B^{*}\end{subarray}}\{a+g,\phi(g),-a+\varphi(g)\}\cup\bigcup\limits_{\begin{subarray}{c}g\in B\end{subarray}}(\{e_{1}+g,e_{2}+\phi(g),-e_{1}-e_{2}+\varphi(g)\}\\ &\cup\bigcup\limits_{g\in B}\left\{c+g,d+\phi(g),-c-d+\varphi(g),-c-g,-d-\phi(g),c+d-\varphi(g)\right\}.\end{array}$
@@
$H=\{0,b,-b,c,-c\}\cup\bigcup\limits_{i=1}^{(|H^{*}|-4)/6}\{b_{i},c_{i},-b_{i}-c_{i},-b_{i},-c_{i},b_{i}+c_{i}\}.$
@@
$\Gamma^{*}=\{(0,b),(0,-b)\}\cup\bigcup\limits_{a\in L^{*}}\{(a,0),(\phi(a),b),(\varphi(a),-b)\}\cup\bigcup\limits_{a\in L}\{(a,c),(-a,-c)\}\cup W.$
@@
$\Gamma^{*}=M^{*}\cup\bigcup\limits_{a\in L}\{(a,b),(-a,-b)\}\cup\bigcup\limits_{a\in L}\{(a,c),(-a,-c)\}\cup W.$
@@
$\Gamma^{*}=M^{*}\cup\bigcup\limits_{a\in L}\{(a,b),(-a,-b)\}\cup\bigcup\limits_{a\in L}\{(a,c),(-a,-c)\}.$
@@
$\begin{array}[]{ll}\Gamma^{*}=&B^{*}\cup\bigcup\limits_{\begin{subarray}{c}g\in B\end{subarray}}\{a+g,-a-g\}\cup\bigcup\limits_{\begin{subarray}{c}g\in B\end{subarray}}\{b+g,-b-g\}\\ &\cup\bigcup\limits_{\begin{subarray}{c}g\in B\end{subarray}}\{e_{1}+g,e_{2}+\phi(g),-e_{1}-e_{2}+\varphi(g)\}\\ &\cup\bigcup\limits_{\begin{subarray}{c}g\in B\\ i=1,2\end{subarray}}\left\{c_{i}+g,d_{i}+\phi(g),-c_{i}-d_{i}+\varphi(g),-c_{i}-g,-d_{i}-\phi(g),c_{i}+d_{i}-\varphi(g)\right\},\end{array}$
@@
$\bigcup\limits_{g\in B}\left\{c_{1}+g,d_{1}+\phi(g),-c_{1}-d_{1}+\varphi(g),-c_{1}-g,-d_{1}-\phi(g),c_{1}+d_{1}-\varphi(g)\right\}.$
@@
$\begin{array}[]{ll}\Gamma^{*}=&\{a,-a\}\cup\bigcup\limits_{\begin{subarray}{c}g\in B^{*}\end{subarray}}\{a+g,\phi(g),-a+\varphi(g)\}\cup\bigcup\limits_{\begin{subarray}{c}g\in B\end{subarray}}\{b+g,-b-g\}\\ &\cup\bigcup\limits_{\begin{subarray}{c}g\in B\end{subarray}}\{e_{1}+g,e_{2}+\phi(g),-e_{1}-e_{2}+\varphi(g)\}\\ &\cup\bigcup\limits_{\begin{subarray}{c}g\in B\\ i=1,2\end{subarray}}\left\{c_{i}+g,d_{i}+\phi(g),-c_{i}-d_{i}+\varphi(g),-c_{i}-g,-d_{i}-\phi(g),c_{i}+d_{i}-\varphi(g)\right\},\end{array}$
@@
Since $|B|=|\Gamma|/20=|L|/4$, $(|L|/2-1,0,0)$ is realizable in $\left(\cup_{g\in B^{*}}\{a+g,\phi(g),-a+\varphi(g)\}\right)$ $\cup\left(\cup_{g\in B}\{e_{1}+g,e_{2}+\phi(g),-e_{1}-e_{2}+\varphi(g)\}\right)$. Since $s<|L|$ and $10\leq s-|L|/2+$ 1, we can find the remaining $s-|L|/2+1\leq|L|/2=2|B|$ zero-sum $3$-subsets in
@@
Let first $e$ be even. Let $m_{1},\dots,m_{|R|/2-1}=2$, $m_{|R|/2},\ldots,m_{|R|/2+|I(\Gamma)|/3-3}=3$, $m_{|R|/2+|I(\Gamma)|/3-2},$ $m_{|R|/2+|I(\Gamma)|/3-1}=4$. Note that $\sum_{i=1}^{|R|/2+|I(\Gamma)|/3-1}m_{i}=|\Gamma^{*}|$. Suppose that there is a subset partition $\{S_{i}\}_{i=1}^{t}$ of $\Gamma^{*}$ such that $|S_{i}|=m_{i}$ and $\sum_{s\in S_{i}}s=0$ for all $i$. Then $R\setminus\left(\bigcup_{i=1}^{|R|/2-1}S_{i}\right)=\{a,-a\}$ for some $a\in R$. Thus $a\in S_{j}$ for some $j\in\{|R|/2,\ldots,|R|/2+|I(\Gamma)|/3-1\}$ but then $\sum_{s\in S_{j}}s\neq 0$ since $I(\Gamma)\cup\{0\}$ is a subgroup of $\Gamma$. Thus for any $\iota_{1},\iota_{2},\iota_{3}\in I(\Gamma)$ there is $\iota_{1}+\iota_{2}\in I(\Gamma)$, $\iota_{1}+\iota_{2}+\iota_{3}\in I(\Gamma)\cup\{0\}$ but $a\notin I(\Gamma)\cup\{0\}$, a contradiction.